# DCA-Bench: A Benchmark for Dataset Curation Agents

### v1

@@ 5 Conclusion
To help the development of LLM agents capable of dataset curation, we present DCA-Bench, a collection of representative dataset issue cases from popular dataset platforms. Instead of fixing predefined issues, DCA-Bench aims to test the agent’s capability to discover hidden data quality issues, a critical initial step in the dataset curation pipeline. To efficiently and effectively evaluate the performance of the Curator agents, we develop an LLM-based Evaluator with carefully designed prompts, which aligns well with human annotators. We conduct a benchmark study on the baseline Curator using DCA-Bench, and the results indicate that while LLMs have potential in real-world dataset curation, further exploration and innovation are needed to fully realize their capabilities.
Limitations
@@
Dataset curation is a complex and comprehensive problem, and the representative cases we collect might not fully cover the entire problem set. Additionally, due to the complexity of dataset files, we cannot guarantee that all the issues in the dataset files in DCA-Bench are labeled. Lastly, we haven’t considered other modality information such as images or audios, which may be helpful to effectively curate multimedia datasets. Based on our work, future studies could explore developing more complex LLM agent systems, best practices for handling multi-modal information in dataset curation, or creating a more realistic simulation environment for testing LLM agents.
Appendix A Appendix
A.1 Real World Dataset Curation Examples
@@
A.2 Human Annotation Codebook
Additional Rules: 1. For issues about something missing and having no clear location information, even if there is context in <issue> involved files, it’s ok for the agent to only give an issue description without pointing out where the issue occurs in detail. 2. You should focus on whether the agent has spotted the issue in ground truth rather than caring about whether the agent includes unrelated examples not present in the context. The inclusion of other irrelevant issues won’t influence its performance in evaluation.
A.3 Issue Type and Tag
@@
To help benchmark users better understand the issue, we assigned each of them several tags. Some tags have a structure of affiliations.
@@
Tab. 6 displays the number of each tag in DCA-Bench in detail.
Table 6: The number of tags in DCA-Bench. Note that the sum of the sub-category might not equal the parent category. For instance, if there’s an ethical or legal risk identified in a dataset document, it should be tagged as [“document-problem”, “ethical/legal-risk”], without including any sub-categories of "document-problem" that are not relevant.
Category Number Sub-category Number
typo 13
wrong-format 6
inappropriate-file 4
ethical/legal-risk 4
internal-discrepancy 11
cross-file-discrepancy 23
data-problem 71 wrong-value 26
missing-value 5
data-leakage 2
apparent-corruption 10
hidden-corruption 25
document-problem 28 wrong-info 17
insufficient-info 8
infrastructure-problem 16 data-access 2
script-code 14
A.4 Calculation of File Number and Context Length
import tiktoken
encoding = tiktoken.encoding_for_model(model)
tokens = encoding.encode(sentence)
token_len = len(tokens)
return token_len
A.5 Strict Evaluation
@@
As defined in Section 3.3, the performance of the Curator is categorized into fail, success, and strict-success. However, accurately distinguishing between success and strict-success in the original triple classification is challenging for Evaluators.
@@
Tab. 7 presents the complete results of the Evaluators’ performance with strict success alignment, compared to human annotations (ground truth). There is a significant drop in the Evaluators’ performance when required to provide a triple classification.
@@
Table 7: The performance of Evaluators with different back-end models, using human annotations as ground truth on 92 (issue, hint) pairs. Compared with binary classification (fail / (success+success+), the triple classification proves to be more difficult, making the evaluation results from the Evaluator less reliable.
@@
GPT-4-0125-preivew 97.83 93.10 100.00 96.43 94.87 89.13 76.52
GPT-4o 92.39 81.25 96.30 88.14 82.59 83.70 66.47
GPT3.5-turbo 68.48 48.21 100.00 65.06 42.15 63.04 38.82
LLama3-70B-Instruct 65.22 45.61 96.30 61.90 36.69 59.78 35.26
A.6 Prompt of Hints Writer

### v2

@@ 6. Conclusion
To help the development of LLM agents capable of dataset curation, we present DCA-Bench, a collection of representative data quality issue cases from popular dataset platforms. Instead of fixing predefined issues, DCA-Bench aims to test the agent’s capability to discover hidden data quality issues, a critical initial step in the dataset curation pipeline. To efficiently and effectively evaluate the performance of the Curator agents, we develop an LLM-based Evaluator with carefully designed prompts, which aligns well with human annotators. We conduct a benchmark study on the baseline Curator using DCA-Bench, and the results indicate that while current state-of-the-art LLMs have great potential in real-world dataset curation, further exploration and innovation are needed to fully realize their capabilities.
@@
Limitations and Future Works. Dataset curation is a complex and comprehensive problem, and test cases we collect might not fully cover the entire problem set. Additionally, due to the complexity of dataset files, we cannot guarantee that all the issues in the dataset files in DCA-Bench are labeled. Lastly, we have not yet explored other modalities such as images or audio, which could enhance the curation of multimedia datasets. Future research could focus on developing more sophisticated LLM agent systems, establishing best practices for managing multi-modal information in dataset curation, or creating realistic simulation environments for testing LLM agents. We hope our work inspires these future studies, and we decide to leave these aspects for further works.
Appendix A Details about examples, statistics, and construction process of DCA-Bench
A.1. Real World Dataset Curation Examples
@@
A.2. Issue Type and Tag
@@
It’s important to distinguish between the total number of files in a test case and the number of files affected by issues. Tab. 5 shows the distribution of issue-affected files in DCA-Bench.
Table 5. Statistics of files involved with issues. On average, each instance has 1.44 files involved in the issue.
Number of Issue-Affected Files Number of Test Cases
1 159
2 54
3 6
$>$ 3 2
To help benchmark users better understand the issue, we further assigned each of them several tags. Some tags have a structure of affiliations.
@@
Example A.1: An issue example reported on Kaggle which involves racial bias
TitleBoston House Prices B feature is RACIST
Meta-Info
Content
Example A.2: An issue example reported on BIG-Bench that involves a discrepancy between dataset files.
TitleMiss aligned static information
Meta-Info
Content
Example A.3: An issue example which has a wrong target label that needs precise factual knowledge to discern
TitleError in 118th Congress data
Meta-Info
Content
Example A.4: An issue example which has a wrong target label that requires translation ability to discern
TitleMistranslation in conlang_translation task?
Meta-Info
Content
A.3. Discussion on Choices of Dataset Platforms and Cases Selections
We mainly focus on two types of dataset platforms. The first type includes HuggingFace and Kaggle, which hosts a huge volume of datasets and offers a dedicated discussion section for individual datasets. The second type is hosted on GitHub, such as BIG-Bench, where the data quality issues are identified from the issue or pull request (PR) sections. Roughly speaking, for HuggingFace and Kaggle, we first select a few popular datasets and then examine the discussion posts relevant to data quality issues. For GitHub-based repositories, we obtain data quality issues by browsing the issue or PR sections.
Our detailed collection process is outlined as follows:
Kaggle:
  1. (1)
    We begin by using Kaggle API to select datasets that have $\geq 5$ discussions with a proper files size $\leq 512$ MB, and collect them into a candidate list.
  2. (2)
    We manually review the discussion sections of these datasets, collecting discussions relevant to dataset issues and compiling them into a table.
  3. (3)
    For each collected discussion, if the dataset issues are confirmed by either the dataset owner or other participants (with at least two reports of the issue), we consider it as verified. If not, we manually verify the issues by checking the corresponding files, which is straightforward due to the dataset’s version history control (in case the issue has been resolved in the current version).
HuggingFace:
  1. (1)
    We follow a similar approach to Kaggle, first selecting candidate datasets that have enough discussions and proper file size by using HuggingFace API
  2. (2)
    Using the same strategy as for Kaggle, we collect discussions relevant to dataset issues and verify them through the same process.
Platforms Hosted on GitHub:
On GitHub, dataset issues are identified from several sources,
  1. (1)
    Issues raised by platform maintainers during the review of new dataset contribution PRs. These are inherently verified by the maintainers.
  2. (2)
    Issues reported by users in the issues section. If these issues are confirmed by the maintainers, they are considered verified (And it’s usually linked to a PR so as to fix it). If not, we manually check the corresponding files ourselves.
We have also surveyed on other dataset hosting platforms, which we found are not ideal for constructing DCA-Bench. Zenodo ([n. d.]) is a platform for safely storing and sharing research data, which fully follows FAIR principles (Wilkinson et al., 2016). However, to our knowledge, it doesn’t support a feedback mechanism, which makes it unclear whether there are issues within those dataset files and whether the descriptions align with the datasets. Compared to HuggingFace and Kaggle, this lack of feedback makes the platform less informative about the issues that concern dataset users, which is why we did not collect datasets from it. Harvard-Dataverse ([n. d.]) faces similar issues, as it has a GitHub repository but does not use it for reporting dataset problems. Domain-specific datasets like the Materials-Data-Facility ([n. d.]) or PANGAEA ([n. d.]) also share the same problem, making it difficult for us to collect issue samples from them, which is why they were less interesting for us. Open-Science-Framework ([n. d.]) seems not to be a standard dataset hosting platform—it is a platform for hosting projects, where the files do not necessarily have to be datasets, so we dismissed further discussions.
Appendix B Details about DCA-Bench Evaluator, Hints Writer and Curator
B.1. Design Details of the Evaluator
In this part, we elaborate on the prompt design of our Evaluator. While it’s natural to ask LLMs to directly classify the Curators’ performance into fail, success, success+, this strategy is poorly aligned with the annotation of human experts as we tested. In addition to common prompt methods (e.g., COT (Wei et al., 2023), role-playing (Kong et al., 2024), delimiter (OpenAI, 2024c), we believe more prompt engineering is needed to improve the performance of the Evaluator, and here is our design:
We chose to have the Evaluator rate on specific metrics because this strategy allows for iterative refinement of the prompt by adjusting the weight and threshold of each metric. This approach decreases uncertainty by introducing fixed numerical comparisons, which provides a clearer direction for tuning, as opposed to relying solely on textual standards, which can be more ambiguous.
In designing the Evaluator’s prompt, we started by selecting a small subset of human-labeled samples. We began with an initial weight distribution of ($0.7,0.2,0.1$) to reflect the relative importance of each metric in Section 2.4. The high initial weight for ”Precise Contextual Evidence” was chosen to align with our focus on whether the Curator identifies annotated issues by providing contextual evidence, as discussed in Section 2.4. We then fine-tuned these weights using the selected subset to better align the Evaluator’s prompt with human judgments.
The motivation behind this design is to make iterative refinement of prompts more effective. Developing a good prompt requires testing it on a subset of human-labeled issues with corresponding outcomes from Curators and then making adjustments based on the outcomes. Unlike the model training process, which benefits from clear gradients and systematic optimization, tuning text prompts is often an empirical process necessitating numerous trials without the guarantee of improvement in its performance. Employing numerical weights and thresholds could enable us to tune the prompt more fine-grained, only changing the weights or threshold based on the results, which share similarities with linear regression, resulting in a more efficient prompt tuning process and better performance. By ”training” on a this human-labeled subset, we finalized weights in the prompt as $0.85$, $0.15$, and $0.05$.
It’s important to highlight that we collected samples for the test set used in Section 3.1 after finalizing the prompt design, ensuring that there is no test data leakage to the prompt tuning.
<rules>
</rules>
The <metrics> are as follows:
"m1": {
"criteria": "Precise Contextual Evidence:
"weight": 0.8,
"range": [0, 1]
"m2": {
"criteria": "Detailed Issue Analysis:
2. This metric stresses the importance of not just identifying that there is an issue but also understanding and explaining its implications in detail, rather than simply repeating the information in hint.",
"weight": 0.15,
"range": [0, 1]
"m3": {
"criteria": "Relevance of Reasoning:
1. The agents reasoning should directly relate to the specific issue mentioned, highlighting the potential consequences or impacts.
2. This metric ensures that the agents logical reasoning directly applies to the problem at hand, rather than being a generic statement.",
"weight": 0.05,
"range": [0, 1]
</metrics>
Reference:
<issue>
<ISSUE>
</issue>
<tips-for-involved-field>
- In the <issue> "involved" field, the content of each file shows lines beginning with "+" or "-" to indicate lines that are newly added or removed to fix the issue. This means the file accessed by the agent is the previous version of the file.
- For issues that only involve adding lines, such issues often have a wider and more general context. Therefore, we should be less strict when evaluating m1, and give it a higher rate.
- For issues that only involve removing lines, such issues often have a narrower and more specific context. Therefore, we should be more strict when evaluating m1.
</tips-for-involved-field>
<hint>
<HINT>
</hint>
<answer>
<ANSWER>
</answer>
response below,
5. The agent is not tasked with solving the problem; its only asked to identify the issue. Therefore, its acceptable for the agent to merely describe the existence of the issue without providing a solution, as long as it is mentioned in the reference involved context.
B.2. The Settings of Evaluators
When developing Evaluators, we test following different models as backends: GPT-4o-2024-1120, GPT-4o-2025-0513 (OpenAI et al., 2024), o3-mini-2025-01-31 (OpenAI, 2024b), GPT-4-0125 preview (OpenAI, 2024a), GPT-3.5-Turbo (OpenAI, 2023b), Llama3.3-70B-Instruct (AI, 2024), Llama3-70B-Instruct (Dubey et al., 2024) Deepseek R1 (DeepSeek-AI et al., 2025a), and Deepseek V3 (DeepSeek-AI et al., 2025b). For OpenAI series of models, we applied OpenAI official API [^6] https://platform.openai.com/docs/api-reference, and set their temperature to $0.0$ to maximize reproducibility and stability. For other open source models, we used SGLang framework (Zheng et al., 2024) for efficient deployment and inference. We didn’t apply qunatizations in order to keep model’s original performance. The dtype is Bfloat16, which is default value. We set temperature to $0.6$ for DeepSeek-R1, following their model cards [^7] https://huggingface.co/deepseek-ai/DeepSeek-R1, while setting temperature to $0.0$ for all other models.
B.3. The Settings of Curators
We implemented several baseline Curators using SGLang (Zheng et al., 2024), LangChain (Chase, 2022), and the OpenAI Assistant API. For open-source models, including DeepSeek-R1 (DeepSeek-AI et al., 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2025b), we deployed them locally using the SGLang framework.
To implement the Retrieval-Augmented Generation (RAG) pipeline, we utilized the LangChain framework. We first built a knowledge vector store using the FAISS (Douze et al., 2024) package and then constructed the retrieval chain using the create_stuff_documents_chain and create_retrieval_chain functions provided by LangChain. To integrate SGLang-deployed models within LangChain, we implemented a custom class named SGLangLLM, following LangChain’s model interface.
Regarding temperature settings, we set the temperature to $1.0$ for OpenAI models and DeepSeek-V3. For DeepSeek-R1, we followed its model card recommendations and set the temperature to $0.6$.
B.4. Human Annotation Codebook for Evaluating Curator’s Performance
Additional Rules: 1. For issues about something missing and having no clear location information, even if there is context in ¡issue¿ involved files, it’s ok for the agent to only give an issue description without pointing out where the issue occurs in detail. 2. You should focus on whether the agent has spotted the issue in ground truth rather than caring about whether the agent includes unrelated examples not present in the context. The inclusion of other irrelevant issues won’t influence its performance in evaluation.
B.5. Additional Results from Human-Alignment Experiment on Evaluator
@@
As defined in Section 2.3, the performance of the Curator is categorized into fail, success, and strict-success. However, accurately distinguishing between success and success+ in the original triple classification is challenging for Evaluators.
@@
Tab. 6 presents the complete results of the Evaluators’ performance with success+ alignment, compared to human annotations (ground truth). There is a significant drop in the Evaluators’ performance when required to provide a triple classification.
@@
Table 6. The performance of Evaluators with different backend models (Binary and Triple Classification), using human annotations as ground truth. We treat succ as the positive class when calculating Precision, Recall, and F1 Score (Binary). Color Legend:   (95–100%),   (85–95%),   (60–85%),   (0–60%). Compared with binary classification (fail / (success+success+)), the triple classification proves to be more difficult, making the evaluation results from the Evaluator less reliable, we therefore leave that part for future works.
@@
gpt-4o-2024-11-20 97.83 100.00 92.59 96.15 94.64 88.04 72.98
gpt-4-0125-preview 96.74 92.86 96.30 94.55 92.22 88.04 73.77
gpt-4o-2024-05-13 92.39 81.25 96.30 88.14 82.59 83.70 66.46
DeepSeek-R1 92.39 81.25 96.30 88.14 82.59 86.96 72.61
DeepSeek-V3 93.48 88.89 88.89 88.89 84.27 83.70 63.99
gpt-3.5-turbo 68.48 48.21 100.00 65.06 42.15 63.04 38.82
Meta-Llama-3-70B-Instruct 69.57 49.09 100.00 65.85 43.68 64.13 41.30
Meta-Llama-3.3-70B-Instruct 88.04 72.22 96.30 82.54 73.73 80.43 61.52
o3-mini-2025-01-31 91.30 100.00 70.37 82.61 77.04 82.61 56.96
B.6. Consistent Evaluation
To evaluate the consistency of ratings produced by this prompting strategy, we ran the Evaluator, backed by GPT-4-0125-preview and GPT-4o-2024-11-20, separately three times on the same outputs generated by the Baseline Curator (GPT-4-0125-preview) using a subset of DCA-Bench (92 data-hint combinations). For GPT-4-0125-preview, the standard deviation is ±$\pm 0.89\%$, while for GPT-4o-2024-11-20, the standard deviation is ±$\pm 0.51\%$, indicating high stability. The detailed numbers are included in Tab. 7.
Table 7. Success rates and standard deviations of multiple runs with different backend models (in percentage)
Model Trial 1 (%) Trial 2 (%) Trial 3 (%) Std Dev (%)
GPT-4-0125-preview 27.17 26.09 28.26 0.89
GPT-4o-2024-1120 27.17 26.09 26.09 0.51
B.7. Negligible Self-Preference and Length Bias of Evaluator
In this section, we discuss the possibility that LLM Evaluators may be biased towards Curators that are backed by the same LLM model (self-preference), and might be biases toward the length of the response (length bias) (Panickssery et al., 2024; et al., 2023). We design two experiments to examine whether these biases are within acceptable range.
Experiment on Self-Preference Bias We used GPT-3.5-Turbo and GPT-4-0125-preview as the backend models for both the Curators and the Evaluators, leading to four combinations of Curator and Evaluator settings. For each Evaluator, we calculated the Average Odds Difference (AOD) (Majumder et al., 2022), a metric that measures fairness by assessing differences in true positive and false positive rates between two groups. A value of zero indicates perfect fairness, with equal model performance across both groups. Higher AOD values suggest increasing disparity, with the model performing unevenly between the groups. The AOD is defined as:
(1) $\text{AOD}=\frac{1}{2}\left(\left|\text{TPR}_{\text{group1}}-\text{TPR}_{\text{group2}}\right|+\left|\text{FPR}_{\text{group1}}-\text{FPR}_{\text{group2}}\right|\right)$
In this context, the GPT-4-0125-preview Curator is privileged, while the GPT-3.5-Turbo Curator is unprivileged. We evaluated the GPT-3.5-Turbo Evaluator five times and the GPT-4-0125-preview Evaluator three times.
Table 8. True Positive Rate (TPR) Results
Evaluator \ Curator GPT-3.5-Turbo GPT4-0125-preview
GPT-3.5-Turbo $100.00\,(\pm 0.00)\%$ $100.00\,(\pm 0.00)\%$
GPT-4-0125-preview $84.44\,(\pm 3.14)\%$ $91.35\,(\pm 3.49)\%$
Table 9. False Positive Rate (FPR) Results
Evaluator \ Curator GPT-3.5-Turbo GPT4-0125-preview
GPT-3.5-Turbo $68.95\,(\pm 1.29)\%$ $56.78\,(\pm 3.18)\%$
GPT-4-0125-preview $0.00\,(\pm 0.00)\%$ $3.71\,(\pm 2.92)\%$
Using the error propagation formula [^8] https://en.wikipedia.org/wiki/Propagation_of_uncertainty and Eq. 1, we get the AOD for GPT-3.5-Turbo is $6.09\,(\pm 1.72)\%$.We can also get that the AOD for GPT-4-0125-preview, which is $5.31(\pm 2.76)\%$.
Since an $AOD\leq 10\%$ is considered fair (Majumder et al., 2022), the bias remains within an acceptable range.
Furthermore, we observe a great alignment between the GPT-4-0125-preview based Evaluator and human annotations, on both the GPT-3.5-Turbo based Curator and the GPT-4-0125-preview based Curator:
Table 10. Comparison of Metrics between Evaluators and Human Annotations
Metric/Curator GPT-3.5-Turbo GPT-4-0125-preview
Accuracy $97.46\,(\pm 0.51)\%$ $96.38\,(\pm 0.51)\%$
Precision $100.00\,(\pm 0.00)\%$ $96.29\,(\pm 2.92)\%$
Recall $84.44\,(\pm 3.14)\%$ $91.36\,(\pm 3.49)\%$
F1 Score $91.53\,(\pm 1.87)\%$ $93.66\,(\pm 0.97)\%$
$\kappa$ value $90.06\,(\pm 2.16)\%$ $91.13\,(\pm 1.31)\%$
These results suggest that using the same model for dataset curation and evaluation does not introduce significant bias in our study, which is acceptable.
Experiment on Length Bias We also conducted an experiment to further investigate this. After the Curator generated the results, we fed its response to another LLM (gpt-4o-2024-08-06) to expand upon its generation. While the most common bias is the preference towards longer responses, we also tested the case where the response is more succinct. The corresponding prompts are listed as follows.
Rephrase to verbose:
You are provided with a description text, and your task is to make it more detailed and elaborate. Ensure that you do not add any new information or content that was not present in the original context. Your goal is to expand upon the existing meaning without altering it. Additionally, maintain the same format as the original answer. Dont mention this prompt in your response.
=================== Start:
Rephrase to succinct:
You are provided with a description text. Your task is to rephrase it concisely while ensuring all essential information is retained. Do not add new information or omit any key details from the original context. Remember to maintain and keep the same format as the original answer. When answer is already short, you dont have to condense it too much. Do not mention this prompt in your response.
=================== Start:
Both prompts are designed to ensure that the only difference between the responses is the length. We used GPT-3.5-Turbo (0125) and GPT-4o-mini (0718) as the underlying models for the Curator and compared the results with responses of different length to see if there was a significant change in the success rate. The experiment was conducted on 92 cases (at the hint level).
Table 11. Average response length of succinct, normal, verbose settings.
Model Name/Running Type Succinct Normal Verbose
GPT-3.5-Turbo $204.8$ $339.1$ $520.0$
GPT-4o-mini $361.4$ $521.8$ $672.9$
Table 12. Accuracy of succinct, normal, verbose settings.
Model Name/Running Type Succinct Normal Verbose
GPT-3.5-Turbo $13.04\%$ $14.13\%$ $13.04\%$
GPT-4o-mini $20.65\%$ $20.65\%$ $19.57\%$
As can be seen, the difference of accuracy is at most $\approx 1\%$ for various response lengths, indicating that there is no significant length bias with our Evaluator.
B.8. Prompt Design of Hints Writer